### v1

# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

Abstract
Stellar mass, binary black hole (BBH) mergers dominates the sources of gravitational wave (GW) events so far detected the LIGO/Virgo/KAGRA (LVK) experiment. The origin of these BBHs is unknown, and no electromagnetic (EM) counterpart has been undoubtedly associated to any of such GW events. The thin discs of active galactic nuclei (AGNs) might be viable environments where BBHs can form and at the same time produce observable radiation feedback. This paper presents new physically motivated light-curve (LC) solutions for thermal flares driven by the remnant of a BBH merger within the disc of an AGN. Following previous analyses, we consider that the BBH likely creates an under-density cavity in the disc prior to its coalescence. Depending on the merger conditions, the black hole (BH) remnant can leave the cavity, interact with the unperturbed disc, and drive a transient BH wind. The wind expels disc material that expands above and below the disc plane and we consider the emission of these plasma ejections as the EM counterpart to the merger GW. We model the LC of such eruptions as mini supernovae explosions finding that stellar mass merger remnants can drive distinguishable flares in optical and UV bands, with time lags of $\sim 10-40$ ($100-400$) days after the GW event, lasting for a few days (weeks) when occurring in AGN with central engines of $\sim 10^{6}$ ($10^{7}$) M. The Vera C. Rubin Observatory can potentially detect the optical component of these EM counterparts up to redshifts of $z\sim 0.5-1$, accordingly. The present LC models are applicable to localise flaring AGNs as sources of multi-messenger emission following GW alerts.
preprint: APS/123-QED
I Introduction
The measurement of gravitational waves (GWs) by the LIGO/Virgo/KAGRA (LVK) experiment opens a new window to explore our Universe in an unprecedented way. Among the types of GWs the experiment can detect, those from stellar mass binary black hole (BBH) mergers dominate the measurements during the LVK O1-O4 observational runs [1, 2, 3, 4]. Currently, the astrophysical environments where such BBHs form are not known, and multiple channels have been proposed as plausible origins [5, 6, 7].
Electromagnetic flares accompanying BBH mergers could unveil their astrophysical hosts. The search for such multi-messenger (MM) associations is the subject of current active research involving follow-up campaigns [8, 9, 10, 11, 12], statistical analyses of observed data [11, 13, 14, 15, 16, 17], as well as theoretical efforts to understand the production of detectable electromagnetic counterparts [18, 19, 20, 21, 22, 23, 24, 25]. GWs and radiation from a common origin provide the means of measuring the distance and redshift of the source independently. In this regard, MM events triggered by BBH mergers can also find cosmological applications [26, 27, 28, 29]. Contrary to mergers of compact objects involving neutron stars, BBH mergers do not produce radiation by themselves and require occurring in gas-rich environments to produce significant EM signals.
The central regions of active galactic nuclei (AGNs) possess suitable conditions to cluster stellar mass black holes [30, 31, 32] and to induce BBH pairing and merging [33, 34, 35, 36, 37]. The scenario of mergers in AGN discs appears to naturally explain puzzling features revealed by LVK measurements, such as the anti-correlation of mass ratio and effective spin [14], as well as the production of intermediate mass black holes as merger remnants [38, 33]. Interestingly, AGN thin discs are environments dense enough where BBH mergers can produce significant radiation feedback [34, 20, 22].
A number of optical AGN flares measured by the Zwicky Transient Facility (ZTF) [39] have been investigated as possible EM counterparts to certain BBH GW events of the LVK experiment [11, 13]. These MM associations are currently debated, mostly because the spatial correlation among the different messengers is challenging [40, 41, 15, 16, 17]. Given the current LVK localization volume in the sky, if AGNs indeed host the observed BBH mergers, the possible EM counterpart of a single GW event should be found among hundreds or thousands of known AGNs [28]. An additional challenge in spotting such counterparts is that AGNs typically exhibit intrinsic variability on timescales from days to years [42, 43], and different astrophysical phenomena can lead to flaring and emission enhancement in the AGN emission [44, 45, 46, 47, 48]. Physical MM models for BBH mergers are then crucial to favor or disfavor candidate AGN flares as GW counterparts.
A number of astrophysical scenarios involving BBH mergers in AGNs leading to EM signals have been analysed in previous works, most of them considering gas feedback launched by the merger remnant. EM signals powered by radiation-driven outflows from the merger remnant have been analysed by [49, 21, 22]. The merger remnant could also produce collimated, relativistic jets, and the interaction of such jets with the disc under different conditions has been analysed by [23, 50, 24] predicting different EM signatures.
In this paper, we analytically derive new optical light-curve profiles of thermal flares energised by the remnant of a BBH merger that occurs within the thin disc of an AGN. We follow the scenario proposed by [22], and consider that the merger likely occurs within an under-density cavity created by the binary in the thin disc prior to coalescence [22, 51, 23, 50]. The radiation counterpart is then energised when the kicked remnant penetrates the unperturbed disc and produces hyper-Eddington outflows. Unlike [22], who focuses on the outflow breakout emission, here we consider the long-term emission produced when the outflow drives disc ejection and the latter expands and cools outside the disc, resulting in the production of UV and optical flares (the breakout emission predicted in [22] mainly falls in the soft X-ray bands).
Figure 1: Sketch of the EM counterpart scenario discussed in this work. Upper: Radial cut of an AGN thin disc where the BBH coalescence occurs within a pre-merger cavity. The merger remnant is kicked after coalescence and penetrates the unperturbed disc region. Middle: While travelling within the AGN disc, the remnant experiences a highly super-Eddington inflow of gas, leading to the production of a quasi-isotropic wind that expels the disc material above and below the disc plane. Lower: The plasma ejections produce an electromagnetic flare after expanding enough to allow the escape of internal photons.
This paper is organised as follows. In the next section, we derive an analytical model for the EM counterpart emission. In Section II.3, we discuss the model parameters of interest and present the spectra of the predicted emission. In Section IV, we derive optical light-curve profiles, compare them with the sensitivity of optical instruments. We present a summary and conclusions in Section V.
II Electromagnetic counterpart: analytical considerations
We consider a BBH of total mass $M_{\bullet}$, embedded in, and co-rotating with an AGN thin disc, at a distance $a$ from the central SMBH of mass $M_{\mathrm{S}}$. According to the analysis of [22], in this situation the BBH coalescence would likely occur within an under-density cavity of size scale equivalent to the Hill radius of the BBH system:
$r_{\mathrm{cav}}\approx a\left(\frac{1}{3}\frac{M_{\bullet}}{M_{\mathrm{S}}}\right)^{1/3}.$ (1)
Following such prediction and for analytical convenience, here we consider a BBH coalescence at the mid-plane of a thin disc of semiheight $h_{\mathrm{d}}$, and within a cylindrical cavity of radius $r_{\mathrm{cav}}$, as depicted in Figure 1.
A burst of GWs is released at coalescence, and the merger remnant is born with a mass $\sim M_{\bullet}$ [^1] The mass of the remnant is typically slightly smaller than the mass of the BBH due to the release of gravitational waves at coalescence. Nevertheless, for practical purposes proposed in this work, we do not make a distinction between the masses of the BBH and the remnant. and with a likely gravitational kick velocity $v_{\mathrm{k}}$ in the frame of the coalescence. The initial condition of the coalescence is uncertain, thus we consider the kick direction as random.
Figure 2: Sketch of the geometry and the characteristic lengths analysed in this work. Upper: Radial cut of the AGN disc similar to the upper panel of Figure 1. The red and blue regions represent the volumes of the disc that we assume to be expelled above and below the disc plane. Lower: Edge-on view of the AGN disc at the coalescence site. The white circle represents the pre-merger cavity, and the dashed straight line represents the trajectory of the kicked remnant that penetrates the unperturbed disc. The parabolic-like area represents the pseudo-cometary structure discussed in the text, formed due to the interaction of the remnant-driven wind with the matter of the AGN disc.
The magnitude of the kick velocity can range within $v_{\mathrm{k}}\sim$ [100, 2000] km s-1 [53, 54], which is generally highly supersonic with respect to the disc material. On the other hand, the radial drift velocity of the disc matter is highly subsonic [55], and thus the cavity crossing time of the remnant $r_{\mathrm{cav}}/v_{\mathrm{k}}$ is much shorter than the timescale of cavity replenishment [22]. Therefore, the interaction of the remnant with the unperturbed disc (and thus the possible production of electromagnetic feedback) crucially depends on the kick direction $\theta_{\mathrm{k}}$ with respect to the disc plane. We expect no interaction and hence no radiation counterpart for remnants with $\theta_{\mathrm{k}}\rightarrow\pi/2$. On the other hand, a significant interaction between the remnant and the unperturbed disc is expected for small $\theta_{\mathrm{k}}$ (see Figure 1), which is the case we focus on in this work.
A kicked remnant of mass $M_{\bullet}$ penetrating the disc volume perturbs the disc material due to its gravitational influence. Such interaction can be initially assessed with Bondi-Hoyle-Lyttleton (BHL) theory. In a medium of gas density $\rho_{\mathrm{d}}$, the classical BHL accretion onto the kicked remnant is approximately $\dot{M}_{\mathrm{B}}\sim 4\pi\rho_{\mathrm{d}}(GM_{\bullet})^{2}/v_{\mathrm{k}}^{3}$. Normalising this inflow rate by the Eddington accretion limit of the remnant, gives for typical parameters of the problem:
[…]
where $n_{\mathrm{d}}$ is the disc number density. Under such highly super-Eddington inflow, the remnant is expected to develop radiation-driven winds and/or bipolar collimated jets. Recent general relativistic magnetohydrodynamic (GRMHD) simulations of BHL accretion performed by [56] show the production of Blandford-Znajek jets [57] by spinning black holes (BHs) with a fixed spin direction, when traveling through a medium with plasma beta $\beta\gtrsim 10$. It is not clear, however, what the accretion disc’s $\beta$ is where BBHs likely coalesce. Furthermore, it is not clear whether the BBH coalescence produces a remnant with a fixed spin direction. BHs with precessing spin may launch precessing jets that could globally resemble wide-angle bipolar outflows. Given the aforementioned uncertainties, in the present work we consider that the remnant drives a quasi-spherical transient wind during its trajectory within the AGN disc. We assume that the remnant captures matter from the environment while simultaneously driving outflows. These inflow-outflow processes have been analysed in the context of dynamical friction of travelling black holes (BHs) with feedback [58, 59], and to explain the emission of ultra-luminous X-ray sources as supercritical accretors [60, 61].
Once the remnant enters the disc, it takes a period to capture enough material to ignite the wind, and we estimate such period as the crossing time of the BHL radius, in the source frame $\Delta t^{\prime}_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{3}$. Under this consideration, a necessary condition for the production of a remnant wind is:
$\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{2}}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}},$ (3)
i.e., the remnant’s trajectory within the disc (RHS of Equation 3) must be larger than the remnant’s BHL radius (see Figure 2). In condition 3, $h_{\mathrm{d}}$ is the local disc semi-height, which we parameterise using the Shakura & Sunyaev (SS) disc model [55].
The wind is quenched when the remnant exits the disc (i.e., when the black hole has insufficient material to fuel the outflow), leaving two expansive ejections of plasma above and below the accretion disc as depicted in Figure 1 (bottom panel). In this work, we consider the emission produced by such disc eruptions as the electromagnetic counterpart to the preceding merger gravitational waves (GWs), and we approximate the associated light curves (LCs) as those of explosive blast waves, similar to approaches in [62, 63, 24]. In Subsection II.1, we estimate the amount of matter and energy in the disc ejections, and in Subsection II.2 we assess their emission.
II.1 The energy and mass of the disc ejections
During its passage through the disc, the remnant drives a transient outflow which we refer to as the ”BH wind”. This BH wind expels disc matter that emerges from the disc on the same and opposite sides where the remnant leaves the disc surface (see Figure 1, bottom). These ejections are referred to as the ”upper” and ”lower” eruptions, respectively.
Following [49, 21], we parameterise the kinetic power of the BH wind as
$L_{\mathrm{w}}=\eta_{\mathrm{w}}\dot{M}_{\mathrm{in}}c^{2},$ (4)
where $\eta_{\mathrm{w}}\ll 1$ is taken as an efficiency parameter and $\dot{M}_{\mathrm{in}}$ is taken as a fraction $\eta_{\mathrm{in}}<1$ of the classical BHL accretion rate:
$\dot{M}_{\mathrm{in}}=\eta_{\mathrm{in}}4\pi\rho_{\mathrm{d}}\frac{(GM_{\bullet})^{2}}{(c_{\mathrm{s}}^{2}+v_{\mathrm{k}}^{2})^{3/2}},$ (5)
where $\rho_{\mathrm{d}}$ and $c_{\mathrm{s}}$ are the local density and speed of sound in the disc, which we obtain using the SS disc model.
The wind kinetic power can also be expressed as
$L_{\mathrm{w}}=(\Gamma_{\mathrm{w}}-1)\dot{M}_{\mathrm{w}}c^{2},$ (6)
where $\dot{M}_{\mathrm{w}}$ and $\Gamma_{\mathrm{w}}$ are the mass loss rate and the terminal Lorentz factor of the wind flow, respectively. Following [64], we consider that the remnant grows at the critical accretion rate $\dot{M}_{\mathrm{cr}}\equiv\eta\dot{M}_{\mathrm{Edd}}=L_{\mathrm{Edd}}/c^{2}$ with $\eta=0.1$, and the excess mass inflow is expelled to form the wind. Since the rate of mass inflow is typically highly super-Eddington in the present problem, the wind mass loss rate can then be approximated as $\dot{M}_{\mathrm{w}}\approx\dot{M}_{\mathrm{in}}$. Under this consideration, the wind terminal velocity $v_{\mathrm{w}}$ is obtained by combining Equations (4) and (6), giving:
$\Gamma_{\mathrm{w}}=1+\eta_{\mathrm{w}},$ (7)
and hence
$v_{\mathrm{w}}=c\left(1-\Gamma_{\mathrm{w}}^{-2}\right)^{1/2}\approx c\sqrt{2\eta_{\mathrm{w}}},$ (8)
where the last approximation is valid when $\eta_{\mathrm{w}}\ll 1$.
The total energy injected by the remnant is calculated as $E_{0}=L_{\mathrm{w}}\Delta t^{\prime}_{\mathrm{w}}$, where $\Delta t^{\prime}_{\mathrm{w}}$ is the time period of wind injection (source frame):
$\Delta t_{\mathrm{w}}=\frac{l_{\mathrm{d}}-l_{\mathrm{HL}}}{v_{\mathrm{k}}},$ (9)
where $l_{\mathrm{d}}=h_{\mathrm{d}}/\sin\theta_{\mathrm{k}}-r_{\mathrm{cav}}/\cos\theta_{\mathrm{k}}$ is the length of the BH trajectory within the disc, and $l_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{2}$ is the distance travelled before the wind is ignited (see Figure 1). We assume equipartition of the total wind power for energising the upper and lower ejections of matter, which is a reasonable approximation for small kick angles $\theta_{\mathrm{k}}$. Based on the aforementioned considerations, the total energy of the plasma ejected above (+) and below (-) the disc plane is parameterised as:
$E_{0}=\eta_{\mathrm{w}}\eta_{\mathrm{in}}2\pi\rho_{\mathrm{d}}\frac{(GM_{\bullet}c)^{2}}{v_{\mathrm{k}}^{4}}\left[\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}-\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{2}}\right].$ (10)
We note that in the small $\theta_{\mathrm{k}}$ regime, lower limits for the wind injection period and for the energy stored in each plasma ejection (equations 9 and 10, respectively) can be estimated as (see Figure 2):
$\Delta t^{\prime}_{\mathrm{w,ll}}$ $\sim\frac{h_{\mathrm{d}}}{v_{\mathrm{k}}}\sim 5.721\,\,\mathrm{days}$
×$\times\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)\left(\frac{a/R_{\mathrm{g}}}{10000}\right)\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)\left(\frac{v_{\mathrm{k}}}{300\mathrm{kms}^{-1}}\right)^{-1},$ (11)
and
$E_{\mathrm{0,ll}}$ $\sim\eta_{\mathrm{w}}\eta_{\mathrm{in}}2\pi\rho_{\mathrm{d}}h_{\mathrm{d}}\frac{(GM_{\bullet}c)^{2}}{v_{\mathrm{k}}^{4}}\sim 2.466\times 10^{47}\,\mathrm{erg}$
×$\times\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{2}\left(\frac{v_{\mathrm{k}}}{300\mathrm{kms}^{-1}}\right)^{-4}$
×$\times\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)\left(\frac{a/R_{\mathrm{g}}}{10000}\right)\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)\left(\frac{n_{\mathrm{d}}}{10^{14}\mathrm{cm}^{-3}}\right),$ (12)
respectively.
The disc ejections are comprised of the material from the BH wind and swept-up disc material, as illustrated in Figure 1, upper. The wind material is accreted from the disc itself onto the remnant. Therefore, we estimate the masses of the upper and lower ejections as:
$M_{0}^{\pm}=\rho_{\mathrm{d}}V_{\mathrm{0}}^{\pm},$ (13)
where $\rho_{\mathrm{d}}$ is the density of the disc material surrounding the pre-merger cavity, and $V_{0}^{\pm}$ are the volumes in the disc above and below the trajectory of the remnant. The kick velocity $v_{\mathrm{k}}$ is typically highly supersonic with respect to the local sound speed of the disc, hence we expect the BH wind to form a bow shock structure. Considering the area $A_{\mathrm{b}}$ defined by the bow shock shape on a plane containing the BH trajectory, we obtain the upper volume $V_{0}^{+}$ as that of the ”wedge” with faces of areas $A_{\mathrm{b}}$ and $\cos\theta_{\mathrm{k}}A_{\mathrm{b}}$ (red region in Figure 1, middle). The lower volume $V_{0}^{-}$ is that of a prism with a base of $\cos\theta_{\mathrm{k}}A_{\mathrm{b}}$ plus the upper volume (blue region in Figure 1, middle):
$V_{0}^{\pm}=\cos\theta_{\mathrm{k}}A_{\mathrm{b}}\times\begin{cases}&\frac{1}{2}h_{\mathrm{d}}-\frac{1}{2}\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}\;\;(\mbox{upper}),\\ \\ &\frac{3}{2}h_{\mathrm{d}}+\frac{1}{2}\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}\;\;(\mbox{lower}).\end{cases}$ (14)
The bow shock area is calculated as:
[…]
which is defined by the bow shock form $R_{\mathrm{b}}(\phi)$ and a maximum angle $\phi_{b}$ limiting the bow shock extension. To define $R_{\mathrm{b}}(\phi)$, we adopt the exact solution of [65] and [66] that describes the bow shock of a spherical wind travelling supersonically through a uniform medium:
$R_{\mathrm{b}}(\phi)=R_{\mathrm{b,0}}\csc\phi\sqrt{3(1-\phi\cot\phi)}.$ (16)
This solution is characterised by the stagnation radius $R_{\mathrm{b,0}}$, which in terms of the quantities of our problem is given by:
$R_{\mathrm{b,0}}=\sqrt{\frac{\dot{M}_{\mathrm{in}}v_{\mathrm{w}}}{4\pi\rho_{\mathrm{d}}v_{\mathrm{k}}^{2}}}.$ (17)
Finally, we set the limiting angle $\phi_{\mathrm{b}}$, such that the back side of the bow shock area is tangent to per-merger cavity, as depicted in Figure 2. Under this consideration, $\phi_{\mathrm{b}}$ is found solving the following equation:
$R_{\mathrm{b}}(\phi_{\mathrm{b}})\cos(\pi-\phi_{\mathrm{b}})=\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}-R_{\mathrm{b,0}}.$ (18)
The way we proceeded to approximate the mass of the plasma ejections given by equations (13)-(18) relies on the formation of a cometary structure due to the interaction of the remnant’s wind and the disc material. We note that such approximation is appropriate when the path of the remnant within the disc is larger than the stagnation radius $R_{\mathrm{b,0}}$. Therefore, in addition to condition (3), we restrict the present analysis to solutions satisfying:
$R_{\mathrm{b,0}}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}.$ (19)
From this condition, the limiting angle of integral (15) is restricted within the interval $\phi_{\mathrm{b}}\in(\pi/2,\pi)$.
To obtain an order of magnitude estimation for the upper and lower volumes given by equation 14, we expand the RHS of equation (14) up to first order in $\theta_{\mathrm{k}}$ (which is the limit of interest) and approximate the area $A_{\mathrm{b}}$ by that of a circle of diameter $l_{\mathrm{d}}$ (being $l_{\mathrm{d}}$ the virtual path of the remnant within the intact disc, see Figure 2), giving:
$V_{0}^{+}\approx\frac{\pi}{8}\left(\frac{h_{\mathrm{d}}^{3}}{\theta_{\mathrm{k}}^{2}}-\frac{2h_{\mathrm{d}}^{2}r_{\mathrm{cav}}}{\theta_{\mathrm{k}}}\right),$ (20)
and
$V_{0}^{-}\approx V_{0}^{+}+\frac{2\pi}{8}\frac{h_{\mathrm{d}}^{3}}{\theta_{\mathrm{k}}^{2}},$ (21)
for the upper and lower volumes, respectively. Normalising these volumes by $R_{\mathrm{g}}^{3}$ (where $R_{\mathrm{g}}$ is the gravitational radius of the hosting SMBH), we then estimate:
$\frac{V_{0}^{+}}{R_{\mathrm{g}}^{3}}$ $\approx 1.060\times 10^{6}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{3}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-2}$
$-5.028\times 10^{5}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{2}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)^{-1/3}$
×$\times\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{1/3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-1},$ (22)
for the upper expelled volume, and
$\frac{V_{0}^{-}}{R_{\mathrm{g}}^{3}}\approx\frac{V_{0}^{+}}{R^{3}_{\mathrm{g}}}+2.121\times 10^{6}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{3}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-2},$ (23)
for the lower volume.
II.2 The emission of the disc ejections
Following [67, 62, 24], we approximate the disc ejections as spherical blast waves in homologous expansion [68, 69, 70]. The initial radii of such spheres $R_{0}^{\pm}$ are set so that they contain the volumes of the ejected material before being energised by the remnant (see quation 14):
$R^{\pm}_{0}=\left(\frac{3V_{0}^{\pm}}{4\pi}\right)^{1/3},$ (24)
where the superscript “$\pm$” labels quantities corresponding to the upper (+) and lower (-) ejections. Since the gas density drastically drops outside the disc, we the external radii of the spheres undergoe free expansions:
$R^{\pm}(t^{\prime})=R^{\pm}_{0}+u_{0}^{\pm}(t^{\prime}-t^{\prime}_{0}),$ (25)
where $u_{0}^{\pm}$ is the expansion velocity and $t^{\prime}$ is the elapsed time in the AGN’s frame, being $t^{\prime}=0$ the moment of the coalescence. The radii given by equation 25 is defined for $t^{\prime}\geq t_{0}$, and we estimate the initial expansion time as:
$t^{\prime}_{0}=\Delta t^{\prime}_{\mathrm{cav}}+\Delta t^{\prime}_{\mathrm{HL}}+\Delta t^{\prime}_{\mathrm{h}},$ (26)
where $\Delta t^{\prime}_{\mathrm{cav}}=r_{\mathrm{cav}}/(v_{\mathrm{k}}\cos{\theta_{\mathrm{k}}})$ is the period in which the remnant enters the unperturbed disc, $\Delta t^{\prime}_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{3}$ is the elapsed time before the BH wind is ignited, and $\Delta t^{\prime}_{\mathrm{h}}=\Delta t^{\prime}_{\mathrm{w}}/2$ is the time in which the idealised spheres are formed, which we take as an intermediate point within the wind injection period (see equation 9). To obtain the expansion velocity, we assume equipartition of the total energy of the ejection (see equation 10) among thermal and kinetic energy, i.e., $E^{\pm}_{\mathrm{0,th}}=E^{\pm}_{\mathrm{0,kin}}=E_{0}/2$. Thus, the expansion velocity is taken as:
$u_{0}^{\pm}=\left(\frac{E_{0}}{M_{0}^{\pm}}\right)^{1/2},$ (27)
where $E_{0}$ and $M_{0}^{\pm}$ are given by equations (10) and (13), respectively.
As we show in Appendix A, the gas of the plasma ejections is initially radiation pressure dominated. In this case, the equation of state is approximated as $P_{0}=\frac{1}{3}a_{\mathrm{r}}T_{0}^{4}$, with $a_{\mathrm{r}}$ being the radiation constant and $T_{0}$ the initial temperature. On the other hand, assuming the idealised situation in which the plasma ejections have a homogeneous distribution of matter, their thermal energy and gas pressure can be related as $P_{0}V_{0}=E_{\mathrm{0,th}}(\gamma_{a}-1)$, with $\gamma_{a}=4/3$ being the adiabatic index appropriate for a radiation-dominated gas. Then, the initial temperature of the plasma ejections is estimated as:
$T_{0}^{\pm}=\left(\frac{E_{\mathrm{0}}}{2a_{\mathrm{r}}V_{0}^{\pm}}\right)^{1/4}.$ (28)
To obtain an order of magnitude estimation for typical values of the initial temperatures $T_{0}^{\pm}$, we use the energy and initial volume of the plasma ejection given by equations (12) and (20), keeping the first term (the dominant one) in the latter. Then we obtain:
$T_{0}^{+}\approx\frac{1}{v_{\mathrm{k}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{1/2}\left(\frac{8\eta_{\mathrm{w}}\eta_{\mathrm{in}}\rho_{\mathrm{d}}}{a_{\mathrm{r}}}\right)^{1/4}=1.062\times 10^{5}\mathrm{K}$
×$\times\left(\frac{300\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)\left(\frac{0.01}{h_{\mathrm{d}}/a}\right)^{1/2}\left(\frac{10000}{a/R_{\mathrm{g}}}\right)^{1/2}\left(\frac{10^{6}\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{1/2}$
×$\times\left(\frac{\theta_{k}}{0.1}\right)^{1/2}\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{1/2}\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)^{1/4}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{1/4}\left(\frac{n_{\mathrm{i}}}{10^{14}}\right)^{1/4},$ (29)
and $T_{0}^{-}\approx(1/3)^{1/4}T_{0}^{+}$. Since equation (12) underestimates $E_{0}$ and equation (20) over estimates $V_{0}$, the estimated temperature given by (29) is an underestimation for $T_{0}^{\pm}$ when using the full expressions for $E_{0}$ and $V_{0}$ (given by equations 10 and 14, respectively).
For simplicity, we do not calculate the rise of the emission and we consider that photons are trapped until their diffusion timescale meets the expansion dynamical time, which occurs in the period [69, 70]:
$\Delta t^{\prime\pm}_{\mathrm{ph}}=\sqrt{\frac{9\kappa_{\mathrm{T}}M_{0}^{\pm}}{4\pi^{3}cu_{0}^{\pm}}},$ (30)
after the expansion is initiated, being $\kappa_{\mathrm{T}}$ the Thompson’s cross section for electron scattering. The bolometric luminosity of the plasma ejection is then calculated as [68, 69, 70]:
$L^{\pm}(t^{\prime})=\begin{cases}L_{0}^{\pm}\exp\left\{-\frac{t^{\prime 2}}{{t^{\pm}_{g}}^{2}}-\frac{2R^{\pm}_{0}t^{\prime}}{u^{\pm}_{0}{t^{\pm}_{g}}^{2}}\right\}&\;\mathrm{for}\;t^{\prime}\geq t^{\prime}_{0}+\Delta t^{\prime\pm}_{\mathrm{ph}},\\ \\ 0,&\;\mathrm{for}\;t^{\prime}<t^{\prime}_{0}+\Delta t^{\prime\pm}_{\mathrm{ph}},\end{cases}$ (31)
being
$L_{0}^{\pm}=\frac{a_{\mathrm{r}}c(2\pi T^{\pm}_{0}R^{\pm}_{0})^{4}}{27\kappa_{\mathrm{T}}M_{0}^{\pm}},$ (32)
the maximum bolometric luminosity, $t^{\pm}_{\mathrm{g}}=\sqrt{2}\Delta t^{\prime\pm}_{\mathrm{ph}}$, and $a_{\mathrm{r}}$ the radiation constant.
The observed spectrum is calculated as black body emission of a spherical volume with the effective temperature
$T^{\pm}_{\mathrm{eff}}(t^{\prime})=\left[\frac{L^{\pm}(t^{\prime})}{4\pi\sigma_{\mathrm{SB}}R^{\pm}(t^{\prime})^{2}}\right]^{1/4},$ (33)
where $\sigma_{\mathrm{SB}}$ is the Stefan-Boltzmann constant. Thus, the flux density of the plasma ejections at the time $t$ and frequency $\nu$ in the observer’s frame (at Earth) is:
$\nu F^{\pm}_{\nu}(t)=\nu^{\prime}\pi\left[\frac{R^{\pm}(t^{\prime})}{D_{\mathrm{L}}}\right]^{2}B_{\nu^{\prime}}\left[T^{\pm}_{\mathrm{eff}}(t^{\prime})\right],$ (34)
where $z$ and $D_{\mathrm{L}}$ are the redshift and distance of the source, $t^{\prime}=t/(1+z)$ and $\nu^{\prime}=(1+z)\nu$ are the time and radiation frequency in the AGN’s frame, and $B_{\nu^{\prime}}$ is the black body spectral radiance.
In this approach, the time delay in the observer’s frame between the GW event and the appearance of the flare is then:
$\Delta t^{\pm}_{\ell}=(1+z)\left(t^{\prime}_{0}+\Delta t^{\prime\pm}_{\mathrm{ph}}\right),$ (35)
where $t_{0}$ is the starting time of the ejection’s expansion in the AGN frame estimated by equation (26), and the period $\Delta t^{\prime\pm}_{\mathrm{ph}}$ is given by equation (30). In the observer’s frame, the bolometric luminosity decays in the characteristic period (see equation 31):
$\Delta t^{\pm}_{\mathrm{d}}=(1+z)\sqrt{\frac{9\kappa_{\mathrm{T}}M_{0}^{\pm}}{2\pi^{3}cu_{0}^{\pm}}}.$ (36)
which we consider as the duration of the flare. Figure 3 presents a pictorial illustration for the period components of counterpart flare discussed in this section.
Figure 3: Schematic representation of the sequence of periods (not to scale) that encompasses the flare counterpart to a BBH merger in the LC model discussed here (solid curve). The $y$-axis is the bolometric luminosity of the LC, with $L_{0}$ being the maximum bolometric luminosity of the BBH-driven flare (see equation 32). The $x$-axis is the elapsed time after the GW event in the source frame. The dashed curve represents the rise of the flare, not modelled in the present approach (see the text).
II.3 The model parameter space
The EM counterparts discussed in this section are attributed to ejections of plasma driven by a kicked remnant of a BBH merger within an AGN disc. We approximate the thermal emission of such disc eruptions as that of spherical blast waves in homologous expansion [68, 70]. The observed emission corresponds to the photosphere flux given by equation 34, which in the formulation presented here is evaluated by specifying: the distance of the source $D_{\mathrm{L}}$, the mass $M_{\mathrm{S}}$ and accretion rate $\dot{M}_{\mathrm{S}}$ of the AGN’s central engine, the radial distance of the coalescence $a$, the mass of the merger remnant $M_{\bullet}$, the angle $\theta_{\mathrm{k}}$ and velocity $v_{\mathrm{k}}$ of the post-merger gravitational kick, and the inflow and outflow efficiencies $\eta_{\mathrm{in}}$ and $\eta_{\mathrm{w}}$, respectively, of the BH remnant. In what follows, we suggest ranges for the aforementioned parameters, suitable for the assumptions on which the present emission model is built.
The source distance of interest ranges within $D_{\mathrm{L}}\sim[300,3000]$ Mpc, which is compatible with the distances of the BBH coalescences observed during the first three runs of Advanced LIGO [71] and with the distances of AGNs potentially observed with current and forthcoming telescopes. We consider SMBHs with masses within the range of $M_{\mathrm{S}}\sim 5\times 10^{5-7}$ M, as such SMBHs dominate the mass function for AGNs in the local Universe [72, 73]. We are interested in EM counterparts with luminosities comparable to or larger than that of the hosting AGN. Such emissions could occur more effectively in thin discs with relatively low accretion rates, and thus we consider systems with $\dot{m}_{\mathrm{S}}\equiv\dot{M}_{\mathrm{S}}/\dot{M}_{\mathrm{S,Edd}}\sim[0.01,0.1]$. Below $\dot{m}_{\mathrm{S}}\sim 0.01$, the disc experiences a transition to a geometrically thick accretion flow [74], where the present analysis does not apply. The location in the disc where mergers are likely to occur is not well constrained; thus, we consider merger locations ranging within $a\sim[10^{3},5\times 10^{4}]R_{\mathrm{g}}$, consistent with results from previous analyses [35, 75]. The remnant masses of interest fall within $M_{\bullet}\sim[20,80]$ M, which encompasses the bulk of BBH merger remnants seen by previous LVK detections [71]. According to the present analysis, the post-merger kick angle should be small enough to allow sufficient interaction of the remnant with the unperturbed disc region. Thus, we suggest kick angles within $\theta_{\mathrm{k}}\sim[3^{\circ},10^{\circ}]$. The kick velocities in the frame of the coalescence are expected to range within $v_{\mathrm{k}}\sim[100,1000]$ km s-1 [76, 53, 54]. Intuitively, we expect the inflow rate of material onto the kicked BH to be lower than the classical BHL accretion rate (corresponding to the case with no outflow) since BH outflows would prevent some gas inflow. Thus, we suggest inflow efficiencies within $\eta_{\mathrm{in}}\sim[0.05,0.5]$. The outcome of hyper-Eddington inflows onto black holes is poorly studied. Conservatively, here we consider a wind efficiency ranging within $\eta_{\mathrm{w}}\sim[0.001,0.01]$, consistent with previous studies on BH super winds [49, 61]. In addition, a given parameter configuration within the intervals discussed above must fulfil the conditions (3) and (19), i.e.,
$\bm{\max}\left\{R_{\mathrm{b,0}},R_{\mathrm{HL}}\right\}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}.$ (37)
to produce valid emission solutions within the present approach.
When employing the present emission model to interpret flares of a particular AGN considering a given GW event, the quantities $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, and $M_{\bullet}$ can be constrained by observations. In this case, the possible predicted flares have five free parameters, namely, $a$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, $\eta_{\mathrm{in}}$, and $\eta_{\mathrm{w}}$.
In the next section, we illustrate spectral energy distributions and LC profiles using parameter configurations within the aforementioned ranges. Such calculations are obtained employing the full mathematical expressions presented in this section; we do not use the approximated expressions, which were aimed at illustrating order of magnitude estimations.
III Emission profiles
The flares discussed here can potentially be identified as EM counterparts provided their emission is comparable to or brighter than that of the hosting AGN. This comparison can be done at a given observed frequency $\nu$ through the ratio:
$r_{\nu}\equiv\frac{F_{\nu}}{F_{\nu,\mathrm{host}}}=\frac{4\pi D_{\mathrm{L}}^{2}\nu^{\prime}F_{\nu}^{\prime}}{\nu L_{\nu,\mathrm{host}}},$ (38)
where $F_{\nu}$ and $F_{\nu,\mathrm{host}}$ are the emission fluxes of the disc eruption (equation 34) and of the hosting AGN, respectively. For simplicity, we consider the host emission as stationary and assess its differential luminosity as a function of the mass $M_{\mathrm{S}}$ and accretion rate $\dot{M}_{\mathrm{S}}$ onto the central SMBH, as done in [23, 24]:
$\nu L_{\nu,\mathrm{host}}=\left(\frac{f_{\mathrm{n}}}{l_{\mathrm{ref}}}\right)\nu L_{\nu,\mathrm{ref}}.$ (39)
In this approach, the reference differential luminosity $\nu L_{\nu,\mathrm{ref}}$ is taken from the average AGN spectrum derived in [77] (their Figure 7), $l_{\mathrm{ref}}$ is the reference luminosity of this spectrum at $\lambda_{\mathrm{ref}}=4400$, and $f_{\mathrm{n}}$ is a modulating factor
$f_{\mathrm{n}}=10^{44}\mathrm{erg}\,\mathrm{s}^{-1}\left(\frac{M_{\mathrm{s}}}{10^{8}\mathrm{M}_{\odot}}\right)\left(\frac{\dot{M}_{\mathrm{s}}c^{2}}{L_{\mathrm{Edd}}(M_{\mathrm{s}})}\right)\left(\frac{10}{f_{\mathrm{c}}}\right).$ (40)
Setting $f_{\mathrm{c}}=3$, we obtain an AGN spectrum consistent with the emission of the corresponding SS thin disc (see Figures 4-5).
We derive LC profiles of the EM counterparts as the sum of the host (equation 39) and the disc eruption (equation 34):
$F^{\pm}_{\nu,\mathrm{flare}}=(1+z)\left(\frac{\nu L_{\nu,\mathrm{host}}}{\nu^{\prime}4\pi D_{\mathrm{L}}^{2}}+F^{\pm}_{\nu^{\prime}}\right),$ (41)
where $\nu^{\prime}$ and $\nu=\nu^{\prime}/(1+z)$ are the emission frequencies at the source and observer’s frames, respectively, and $z$ is the source redshift which we relate to the luminosity distance $D_{\mathrm{L}}$ through a $\Lambda$CDM cosmological model constrained with Planck 2018 results [78].
Non-blazar AGNs typically exhibit stochastic variability in optical and X-ray bands representing about 10% of their average emission over time scales ranging from days to years [42, 43]. The origin of such variability is not established, as it could be driven by multiple processes in the AGN disc. Therefore, here we consider $r_{\mathrm{th}}=0.25$ as a reasonable threshold to distinguish the flares predicted in this work from their host emission, and define as detectable emission solutions leading to $r_{\nu}\geq r_{\mathrm{th}}$ (see equation 38).
In the following subsections, we illustrate spectral energy distributions (SEDs) and LC profiles predicted by the present model using a fiducial BH remnant of $40$ M with inflow/outflow efficiencies of $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.005$. The other model parameters are varied as indicated.
III.1 Spectral energy distributions
The EM counterpart discussed here is approached as blackbody emission from the photosphere of the disc eruption (see equation 34). As such, the appropriate frequency for observing these flares is $\nu_{\mathrm{pk}}\sim k_{\mathrm{B}}T_{\mathrm{eff}}/[(1+z)h]$ and somewhat lower frequencies. We note that, at maximum bolometric luminosity, the effective temperature $T_{\mathrm{eff}}$ scales inversely with the disc height (see equations 29-33). Therefore, we expect the flares to shine at lower frequencies as the hosting SMBH is larger. This is illustrated in Figures 4-5, where we present SEDs of the disc eruptions considering AGNs with central engines of $10^{6}$ and $10^{7}$ M.
The SEDs displayed in Figure 4 correspond to an “upper eruption” of a merger with a kick angle of $\theta_{\mathrm{k}}=5^{\circ}$, in an AGN with $M_{S}=10^{6}$ M, located at $D_{\mathrm{L}}=1000$ Mpc. In each frame, we display the SEDs of the disc eruption at different observed times after the detection of the GW event, as labelled, starting with the time when photons can escape from the plasma ejection (given by equation 35). We over-plot the spectra of the associated disc emission (obtained with the SS model, thick dashed curve) and of the host galaxy given by equation 39 (thick solid curve). The vertical lines indicate optical ($g$-band), near UV, and far UV frequencies, as labelled. Left/right panels of Figure 4 consider different merger locations $a$ and upper/lower panels different kick velocities $v_{\mathrm{k}}$, as indicated. These values for $a$ and $v_{\mathrm{k}}$ are chosen so that condition 37 is fulfilled.
Figure 4 suggests that merger flares in discs with SMBHs of $\sim 10^{6}$ M have time lags (after the GW event) of week scales and radiate most of their luminosity at far ultraviolet and somewhat higher frequencies. These flares can clearly surpass the host emission at far ultraviolet. Certain solutions, especially those with relatively low kick velocities (near the threshold where condition 37 is satisfied), produce emission comparable to the host at optical bands.
Figure 5 shows SEDs of the merger flare with the same parameter set as in Figure 4, but considering a central SMBH of $10^{7}$ M, and merger locations and kick velocities consistent with condition 37. As expected, these flares peak at lower frequencies than in the $M_{\mathrm{S}}=10^{6}$ M case, and can be suitably detected at optical and UV frequencies. On the other hand, the time scale for the flare lags and durations increases by about an order of magnitude with respect to the $M_{\mathrm{S}}=10^{6}$ M case.
Following the trend exhibited by the SEDs of Figures 4 and 5, one might expect that mergers in AGNs with SMBHs of $10^{8}$ M and heavier can be detectable at optical and infrared frequencies. Nevertheless, the flare time lags in these larger AGNs would be of year scales (we analyse time lag and time duration in more detail in Section III.3), and therefore the multi-messenger associations of real data with such characteristics could be non-trivial. We therefore suggest that the emission mechanism discussed in this paper is suitable for seeking EM counterparts in hosts with $M_{\mathrm{S}}\sim 10^{6-7}$ M.
Figure 4: SED snapshots of the disc eruption (the merger EM counterpart) in an AGN with a central engine of $10^{6}$ M. The SEDs are calculated at different post-merger time steps $t_{\mathrm{pGW}}$ as indicated by the legends in each frame. The emission of the associated thin disc and host galaxy are over-plotted with the thick dashed and solid curves, respectively. The vertical lines indicate the optical g-band, near UV, and far UV frequencies. Curves on the left and right panels are calculated assuming a merger at $9000$ and $20000R_{\mathrm{g}}$ from the central SMBH, whereas upper and lower panels consider kick velocities of 300 and 500 km s-1. Other model parameters shared by all the panels are specified in the text.
Figure 5: Same as in Figure 4, but considering an SMBH mass of $10^{7}$ M, and different merger distances $a$ and kick velocities $v_{\mathrm{k}}$ as indicated.
III.2 Optical light-curves
In the previous subsection, we show that the EM counterpart can be comparable to or brighter than the emission of the host galaxy at optical bands in AGNs with central engines of $10^{6-7}$ M. In this subsection, we derive LC profiles of the counterpart in the optical g-band, illustrating the “upper” and “lower” disc eruptions, the effect of the AGN’s accretion rate, and the remnant’s kick velocity. Furthermore, we consider different redshifts for the source to compare the emission with the limiting magnitude of different optical instruments.
Figure 6 displays the profiles of possible LCs driven by a BBH merger at $5000$ $R_{\mathrm{g}}$ from an SMBH of $10^{6}$ M with a kick angle of $\theta_{\mathrm{k}}=5^{\circ}$ of the merger remnant. Curves in the left and right panels of Figure 6 are obtained assuming an AGN accreting at 0.01 and 0.05 Eddington accretion rates, respectively. In each frame, upper, middle, and lower curve families are obtained assuming different redshifts of the source, as indicated. Curves with different line styles are obtained using different kick velocities as labelled, whereas red and blue curves correspond to the upper and lower disc eruptions. The LCs are calculated as given by equation 41 and converted to AB magnitudes employing the Python package Speclite [^2] https://speclite.readthedocs.io/en/latest/index.html with the $g$ filter of the Sloan Digital Sky Survey (SDSS, Blanton et al. 80). For reference, the horizontal lines on the right in each panel indicate the limiting magnitude of different optical instruments, namely DDOTI (1000s exposure, 3-$\sigma$, white filter [10, 81]), ZTF (30s exposure, 5-$\sigma$ detection, g filter [39]), SPLUS (3$\times$30s, 3-$\sigma$, g filter [82, 83]), DECam (60s exposure, 5-$\sigma$ detection, g filter [28]), and the forthcoming Vera C. Rubin Observatory ($2\times 15$s exposure, 5-$\sigma$ detection, g filter [84]), as labelled.
Figure 6 shows how the flare profiles are impacted by the remnant’s kick velocity $v_{\mathrm{k}}$ and by the SMBH accretion rate $\dot{m}_{\mathrm{S}}$. The emission time lag is shorter for larger values of $v_{\mathrm{k}}$ since the disc crossing time (and hence the time for the production of the disc eruptions) is reduced (see equation 26). Furthermore, the amplitude of the flare is reduced as $v_{\mathrm{k}}$ increases since the energy stored in the disc ejections is approximately proportional to $v_{\mathrm{k}}^{-4}$ (see equation 10). We note that the flux enhancement relative to the (here assumed) stationary AGN emission is smaller as the AGN accretion rate $\dot{m}_{\mathrm{S}}$ is larger. This suggests that EM counterparts can be more distinguishable from their hosts in discs with relatively low accretion rates. On the other hand, for larger $\dot{m}_{\mathrm{S}}$ the overall flux (i.e., AGN plus merger counterpart) increases, hence increasing the distance up to which the source can be detected. These LCs show, for instance, that in AGNs of $10^{6}$ M accreting at $\dot{m}_{\mathrm{S}}=0.01-0.05$, the flares could be detected by ZTF up to $z\sim 0.04-0.09$, and up to $z\sim 0.1-0.3$ by forthcoming Rubin. The curves show that upper eruptions appear always somewhat before and have shorter durations than the lower counterparts. This is because upper ejections are always less massive than the lower ones, and therefore their photon diffusion time (see equations 30 and 27) and flare characteristic period (equation 36) are always shorter in the upper flares.
In Figure 7, we present possible LCs of a flare triggered by a merger with the same parameters as in Figure 6, but considering an SMBH of $10^{7}$ M, and exploring different values for the kick velocity and the redshift of the source, as indicated. The flare time lags and durations are of the order of hundreds and tens of days, respectively, which represents an increase of about an order of magnitude compared to such periods in Figure 6. However, the effect on the LC profiles due to variations of $v_{\mathrm{k}}$ and $\dot{m}_{\mathrm{S}}$ is qualitatively similar. We also note that in this $M_{\mathrm{S}}=10^{7}$ M case, lower values for $v_{\mathrm{k}}$ are allowed by condition 37, and that the overall emission can be detected at larger redshifts compared to the $M_{\mathrm{S}}=10^{6}$ M case. For instance, curves in Figure 7 show that in discs accreting at $\dot{m}_{\mathrm{S}}=0.01-0.05$, the flares could be detected by ZTF up to $z\sim 0.1-0.3$, and up to $z\sim 0.5-1$ by forthcoming Rubin.
The SEDs and LCs presented in Figures 4-7 suggest that, in general, the flares are more distinguishable from their host as the merger distance $a$ and the kick velocity $v_{\mathrm{k}}$ are smaller.
Figure 6: LCs in the optical $g$-band of the EM counterpart model for a BBH merger in an AGN disc. The figure illustrates possible LC profiles driven by a merger with the following model parameters: $M_{\mathrm{S}}=10^{6}$ M, $a=5000R_{\mathrm{g}}$, $M_{\bullet}=40$ M, $\theta_{\mathrm{k}}=5^{\circ}$, $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.005$ (see the text for details). We explore different redshifts for the source which results in the upper, middle, and lower curve families as well as different kick velocities, as labelled. Red and blue curves correspond to the “upper” and “lower” eruptions of the model. Left and right panels correspond to AGNs accreting at different rates $\dot{m}_{\mathrm{S}}$, as indicated. The horizontal lines indicate limiting magnitudes of different optical instruments as labelled (see details in the text).
Figure 7: Same as in Figure 7, but considering an SMBH of $M_{\mathrm{S}}=10^{7}$ M and different kick velocities $v_{\mathrm{k}}$ and source redshifts $z$, as indicated.
III.3 The flare delay and duration.
The SEDs and LCs examples displayed in Figures 4-7, show that the time lag $\Delta t_{\ell}$ and the duration $\Delta t_{\mathrm{d}}$ of the flares are notably impacted by the mass of the hosting SMBH $M_{\mathrm{S}}$, the location of the merger $a$, and the kick velocity $v_{\mathrm{k}}$. Here we describe how the present emission model predicts $\Delta t_{\ell}$ and $\Delta t_{\mathrm{d}}$ in a more systematic way.
In Figure 8, we display $\Delta t_{\ell}$ given by equation 35 as a function of the SMBH mass $M_{\mathrm{S}}$. In this example, we consider a source located at $D_{\mathrm{L}}=1000$ Mpc and derive the $\Delta t_{\ell}$ vs $M_{\mathrm{S}}$ curves exploring different values for $a/R_{\mathrm{g}}$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, and $\dot{m}_{\mathrm{S}}$. We restrict these curves to fluxes with ratios $r_{\nu}\geq 0.25$ in the optical $g$-band (see equation 41) together with condition 37. Thus, the curves in Figure 8 initiate/terminate when the aforementioned conditions are no longer satisfied. In each panel, the curves with different values of $a/R_{\mathrm{g}}$ are labelled with different line styles, whereas curves of different kick velocities are labelled with different colours. Computations in the upper and lower panels are obtained assuming different kick angles $\theta_{\mathrm{k}}$, whereas left and right panels correspond to different SMBH accretion rates $\dot{m}_{\mathrm{S}}$, as labelled.
The curves in Figure 8 show that the time lag of the EM counterpart is approximately correlated to the SMBH mass as $\propto M_{\mathrm{S}}^{\alpha_{\ell}}$ with $\alpha_{\ell}\approx 1$. As already noted in the SEDs and LCs of previous subsections, Figure 8 shows that $\Delta t_{\ell}$ increases for larger merger locations and is diminished for larger kick velocities. The parameter space leading to visible flares is reduced for larger values of $\dot{m}_{\mathrm{S}}$ and $\theta_{\mathrm{k}}$, as also noted in the previous subsections. Interestingly, the condition $r_{\nu}\geq r_{\mathrm{th}}$ together with condition 37 associates the mass of the hosting SMBH with a characteristic kick velocity that would produce detectable flares. For instance, we see that in AGNs with SMBHs of $10^{6}$ M, visible flares would be more likely produced by remnants with kicks of $\sim 350$ km s-1, and of $\sim 150$ km s-1 in discs around SMBHs of $10^{7}$ M. Such $M_{\mathrm{S}}-v_{\mathrm{k}}$ association appears more defined as the SMBH accretion rate is larger.
In Figure 9, we display the flare duration given by equation 36 and associated to the curves of Figure 8. This figure shows that $\Delta t_{\mathrm{d}}$ is generally one order of magnitude smaller than the flare time lag. The $\Delta t_{\mathrm{d}}-M_{\mathrm{S}}$ relation is approximately linear, similar to $\Delta t_{\ell}$. However, we note that the effect of the merger location $a$ is less pronounced in the flare duration than in the flare time lag.
The present EM counterpart model is built such that the flare time lag is composed of a sequence of periods, namely the cavity crossing time $\Delta t_{\mathrm{cav}}$, the time for the outflow formation $\Delta t_{\mathrm{HL}}$, the mean time for the formation of the disc ejection $\Delta t_{\mathrm{h}}$, and the photon escaping time $\Delta t_{\mathrm{ph}}$ (see Section II.2 for details). The contribution of these periods to the flare time lag varies according to the model parameter configuration. Generally, $\Delta t_{\mathrm{cav}}$ and $\Delta t_{\mathrm{h}}$ are the dominant components, as illustrated in Figure 10. In this figure, we display the components of the time lag as a function of merger location, considering an AGN at a distance of 1000 Mpc, accreting at a rate of $\dot{m}_{\mathrm{S}}=0.01$, and within the conditions 37 and $r_{\nu}\geq r_{\mathrm{th}}$ (similarly as in Figures 8-9). Curves in the upper, middle, and lower panels are calculated with different kick angles as indicated. Based on the $M_{\mathrm{S}}-v_{\mathrm{k}}$ relation discussed in the previous subsection, the curves in the left panels are obtained assuming an SMBH of $M_{\mathrm{S}}=10^{6}$ M with a remnant of $v_{\mathrm{k}}=350$ km s-1, whereas curves in the right panels assume $M_{\mathrm{S}}=10^{7}$ M and $v_{\mathrm{k}}=150$ km s-1. Figure 10 shows that the period $\Delta t_{\mathrm{h}}$ dominates in the case of very small kick angles. On the other hand, the contribution of $\Delta t_{\mathrm{cav}}$ rapidly increases as $\theta_{\mathrm{k}}$ is larger, becoming dominant for $\theta_{\mathrm{k}}\gtrsim 8^{\circ}$ (still in the small kick angle regime). The photon escape time $\Delta t_{\mathrm{ph}}$ is generally subdominant within the parameter space of interest.
Figure 8: Time lag of the counterpart flare after the GW event as a function of the hosting SMBH mass. Curves in the left and right panels are obtained considering different SMBH accretion rates, whereas upper and lower panels depict different kick angles for the merger remnant, as labelled. In each panel, different curves correspond to different locations of the merger and different kick velocities of the remnant, as indicated. The curves terminate when conditions 37 and $F_{\nu}/F_{\nu,\mathrm{host}}\leq 0.25$ are no longer satisfied (see the text for further details).
Figure 9: Same as in Figure 8, but displaying the $\Delta t_{\mathrm{d}}$ vs $M_{\mathrm{S}}$ curves.
Figure 10: Examples of the flare time lag and its components as a function of the merger location $a$. These periods are calculated in the observer frame. The curves initiate when condition 37 is fulfilled. The parameter configurations for the curves in each panel are specified in the plot labels and in the text.
IV Optical counterpart candidates to BBH mergers
Graham et al. [11, 13] have proposed a number of optical flares measured by the ZTF from known AGNs as possible EM counterparts associated with LVK BBH merger events. These counterpart candidates are reported with time delays $\Delta t_{\ell}\lesssim 200$ days after the associated GW events, in AGNs situated at $z>0.3$ with central engines of $M_{\mathrm{S}}>10^{8}$ M. Such MM associations are currently under debate [15, 16, 17], mostly because of the large spatial error of the LVK events. Most of these candidate flares are not compatible with the present MM scenario, which predict optical flares with $\Delta t_{\ell}\lesssim 200$ days provided $M_{\mathrm{S}}<10^{7}$ M, and such flares could be detected up to $z\sim 3$ with the sensitivity of the ZTF. In AGNs with $M_{\mathrm{S}}\gtrsim 10^{8}$ M, the flare delays would be likely of year scales under the MM scenario discussed here (see Figure 8). Thus, if the aforementioned candidate flares are real EM counterparts, they would be better explained by mechanisms with faster interaction among the BBH remnant and the AGN disc, such as the re-directed jet scenario discussed in [23, 50].
On the other hand, more precise measurements for the properties of the aforementioned AGN candidates could change the scenario that better explain the potential MM association. For concreteness, we note the mass of the central engine in the AGN J053408.41+085450.6 is currently not accurately measured as pointed out in [13]. Furthermore, J053408.41+085450.6 is reported with photometric redshift, which is a less accurate method compared to spectroscopic redshift measurements of the other AGN candidates. If upgraded measurements lead to an SMBH mass close to $\sim 10^{7}$ M and a redshift $z\lesssim 0.3$ for J053408.41+085450.6, the optical flare peaking at $\sim$ MJD 58900 from this source could be naturally explained by the present emission model as the counterpart to the GW event with ID GW190731_140936. We leave such investigation as the subject of a forthcoming work.
V Summary and conclusions
In this paper, we present a physical, analytic model that predicts the time delay, duration, spectrum, and optical LC of EM feedback from stellar mass BBH mergers within thin AGN discs. Such environments have been proposed as promising hosts for the observed GW events associated to BBH mergers measured by the LIGO/Virgo/Kagra experiment. The EM processes considered come from the thermal emission of plasma eruptions expanding outside the disc, expelled from the disc by the BH merger remnant. In our analytic approach (see section II for details), we assume that the BBH is embedded in the disc and co-rotating. After the merger, the remnant is gravitationally kicked out, accretes from the unperturbed disc and produces a quasi-spherical outflow which expels disc material. These ejections expand and cool outside the AGN disc, producing thermal flares which we model here.
In our formulation, the EM counterpart is determined by nine parameters characterising the hosting AGN ($D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$), the BBH merger ($a$, $M_{\bullet}$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$), and the remnant’s outflow ($\eta_{\mathrm{in}}$, $\eta_{\mathrm{w}}$; see Section II for details). When the model is applied to interpret a particular LC dataset, MM observations can constrain the parameters $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, and $M_{\bullet}$, and then the possible LCs predicted by the model would depend on the other five parameters.
Under the present scenario, BBH mergers produce EM feedback with the following features.
  1. 1.
    The flare emission can be comparable to or outshine that of the hosting AGN at UV-optical frequencies, for mergers occurring in AGNs with SMBHs of $10^{6-7}$ M.
  2. 2.
    The EM counterparts are produced by kicked remnants that leave the AGN disc in the same or opposite direction to the observer (see Figure 1), which we denominate as “upper” and “lower” flares, respectively. The upper flares are somewhat brighter and have slightly shorter time delays than their lower flare counterparts.
  3. 3.
    The kick angle relative to the disc plane $\theta_{\mathrm{k}}$ is a crucial parameter that modulates the flare profile and determines whether EM feedback is possible (see e.g., equations 10 and 37). Within the parameter space discussed here, our results suggest that significant EM counterparts are possible when $\theta_{\mathrm{k}}\lesssim 10^{\circ}$.
  4. 4.
    In AGNs with SMBHs of $\sim 10^{6}(10^{7})$ M, distinguishable flares have time delays of a few tens (hundreds) of days and require remnants with kick velocities $\gtrsim 300$ (100) km s-1. For hosts with SMBHs of $\sim 10^{8}$ M (and heavier), the flares have time lags of thousands of days (and longer). We then suggest that AGNs with SMBHs masses $\lesssim 10^{7}$ M are the most suitable hosts for localising EM counterparts under the present scenario.
  5. 5.
    EM counterparts from AGNs with SMBHs of $10^{6-7}$ M are suitable targets for time-domain UV space telescopes like the forthcoming ULTRASAT [85] and UVEX [86], as well as for optical time domain surveys (see Figures 4-5).
  6. 6.
    We estimate that ZTF can detect the flares predicted here up to $z\sim 0.04-0.09$ from AGNs with $M_{\mathrm{S}}=10^{6}$ M, and up to $z\sim 0.1-0.3$ when $M_{\mathrm{S}}\sim 10^{7}$ M. Rubin could observe the flares up to $z\sim 0.1-0.3$ and $z\sim 0.5-1$ when $M_{\mathrm{S}}\sim 10^{6}$ and $10^{7}$ M, respectively.
The MM scenario discussed here was previously proposed by [22]. These authors focused on the emission produced when the remnant’s outflow breaks out from the disc, predicting X-ray emissions. In contrast to [22], here we focus on the long-term emission of the plasma ejection as it expands outside the disc plane, predicting flares at optical-UV frequencies as described above.
The emission of the disc eruptions discussed here relies on complex physics that we have simplified for analytical purposes. For instance, the morphology of the disc ejections (which are the sources of the EM counterparts) can be far from spherical, and their evolution could be more accurately described by including continuous injection of matter during the early expansion stages when the remnant’s outflow is active. Such corrections could be suitably addressed through magneto-hydrodynamical simulations of the present problem combined with numerical radiative transfer. Additionally, the outcome of highly super-Eddington inflow/outflows in travelling black holes is currently not well studied, and here we modelled such processes by introducing the efficiency parameters $\eta_{\mathrm{in}}$ and $\eta_{\mathrm{w}}$. Thus, if the emission scenario analysed here is indeed correct, MM observations from BBH mergers in AGNs could also serve to probe the physics of hyper-Eddington accretion onto black holes.
Acknowledgements.
JCRR acknowledges support from Rio de Janeiro State Funding Agency FAPERJ, grant E-26/205.635/2022. CdB acknowledges the financial support from CNPq (316072/2021-4) and from FAPERJ (grants 201.456/2022 and 210.330/2022) and the FINEP contract 01.22.0505.00 (ref. 1891/22). RN acknowledges support from CNPq through a Bolsa de Produtividade and NASA through the Fermi Guest Investigator Program (Cycle 16). The authors made use of Sci-Mind servers machines developed by the CBPF AI LAB team and would like to thank P. Russano and M. Portes de Albuquerque for all the support in infrastructure matters.
Appendix A The initial pressure of the plasma ejection
In this appendix, we evaluate whether the disc ejections of plasma are initially gas pressure or radiation pressure dominated. We consider the total pressure as the sum of gas plus radiation pressure:
$P=P_{\mathrm{g}}+P_{\mathrm{r}},$ (42)
where
$P_{g}$ $=2n_{i}k_{B}T,$ (43)
$P_{r}$ $=\frac{1}{3}a_{r}T^{4}.$ (44)
In equation 43, $n_{i}$ denotes the number density of ions, and the factor of two corresponds to a fully ionised gas. To determine which pressure component is dominant, we consider the parameter
$\beta_{r}\equiv\frac{P_{g}}{P},$ (45)
where the cases $\beta_{r}\ll 1$ and $\beta_{r}\sim 1$ correspond to radiation pressure and gas pressure dominated regimes, respectively. In terms of $\beta_{r}$, the pressure components are:
$P_{g}$ $=\beta_{r}P,$ (46)
$P_{r}$ $=(1-\beta_{r})P.$ (47)
Combining equations (42)–(47), the parameter $\beta_{r}$ can be obtained by solving the quartic equation
$A\beta_{r}^{4}+\beta_{r}-1=0,$ (48)
where $A$ is the dimensionless parameter
$A=\frac{a_{r}P^{3}}{3(2n_{i}k_{B})^{4}}.$ (49)
We estimate a lower limit for the typical values of the parameter $A$ using $P\approx E_{\mathrm{0,th}}/V_{0}$, which we underestimate using $V_{0}\approx(3\pi/8)h_{\mathrm{d}}^{3}/\theta_{\mathrm{k}}^{2}$, where $V_{0}$ is an overestimation for the initial volume of the disc eruption (see equation 21), and $E_{\mathrm{0,th}}=E_{\mathrm{0,ll}}/2$ given by equation 12 is an underestimation for $E_{0}$. Then we obtain:
$A_{\mathrm{ll}}$ $\sim\frac{16}{81}\frac{a_{\mathrm{r}}m_{\mathrm{p}}^{3}}{k_{\mathrm{B}}^{4}}\frac{(\eta_{\mathrm{w}}\eta_{\mathrm{in}})^{3}}{n_{\mathrm{i}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{6}\frac{1}{v_{\mathrm{k}}^{12}}$ (50)
$\sim 4.3\times 10^{6}\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)^{3}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{3}\left(\frac{10^{\mathrm{14}}\mathrm{cm}^{-3}}{n_{\mathrm{i}}}\right)$
×$\times\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{6}\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{6}\left(\frac{0.01}{h_{\mathrm{d}}/a}\right)^{6}\left(\frac{9000}{a/R_{\mathrm{g}}}\right)^{6}$
×$\times\left(\frac{10^{6}\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{6}\left(\frac{400\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)^{12},$
Considering values for $A$ within $(10^{4},10^{8})$, solutions to equation 48 give values for $\beta_{\mathrm{r}}$ within $(10^{-5},0.1)$. Therefore, we can safely assume that the plasma eruptions discussed in the paper are initially radiation pressure dominated.